OpenStudy (anonymous):

2 bananas are to be selected from a group of 5. In how many ways can this be done?

OpenStudy (anonymous):

When you pick the first banana, there are 5 to choose from, when you pick the second, there will be four. (5)(4) = 20. But, if i pick banana 1 then banana 3, or if i had picked banana 3 then 1, thats the same outcome, so I need to divide by 2. Answer is 10 i believe.

OpenStudy (anonymous):

thats not an answer option....

OpenStudy (anonymous):

hmmm...what are the answer choices?

OpenStudy (anonymous):

1, 20, 120, and not "here" i get worried iwth the not here's :s

OpenStudy (anonymous):

lol, yeah, those "not here" answers freak me out too. Another way to compute the answer is to do \[\left(\begin{matrix}5 \\ 2\end{matrix}\right)\] which is read "5 choose 2", or, "the number of ways to choose 2 objects at a time from 5 objects" have you seen this before?

OpenStudy (anonymous):

yes, and i guess the more i think about it, 10 seems right......oiy!

OpenStudy (anonymous):

yeah, the answer is still 10 after computing that, I'll put my money on "Not Here" :)

OpenStudy (anonymous):

10

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

here we go

OpenStudy (anonymous):

hey i have another one, ill do it and could u check it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

2 bananas are to be selected from a group of 6. In how many ways can this be done? i think the answer is 30?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

but in that last one we multiplied

OpenStudy (anonymous):

When i pick the first banana, there are 6 to choose from, then 5 for the second. (6)(5) = 30 But (just like the last problem) i have to divide by 2 Answer should be 15

OpenStudy (anonymous):

ohhhhhhh now i seee!

OpenStudy (anonymous):

You own 10 hats and are taking 2 on vacation. In how many ways can you choose 2 hats from the 10? so for this one the answer would be......45?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

general formula for choosing r things out of n things no of ways = n!/ r!(n-r)!

OpenStudy (anonymous):

SWEET! i think im actually understanding this stuff! ooh quick while im on a role lets do another one!

OpenStudy (anonymous):

9 students volunteer for a committee. How many different 7-person committees can be chosen? 181,440????

OpenStudy (anonymous):

That is a good formula to know :) \[\left(\begin{matrix}n \\ r\end{matrix}\right) = \frac{n!}{r!(n-r)!}\]

OpenStudy (anonymous):

or would it be 362,880

OpenStudy (anonymous):

So you want to compute \[\left(\begin{matrix}9 \\ 7\end{matrix}\right) = \frac{9!}{7!(9-7)!}\] = 36

OpenStudy (anonymous):

oh....hmmmmlets try a diff one i will get 1 i have to !!!!

OpenStudy (anonymous):

oops wrong 1

OpenStudy (anonymous):

sure thing, i would really try to get the formula down, for smaller numbers (like the banana and hat problem) you can rationalize an answer. But for larger numbers the formula works best

OpenStudy (anonymous):

6 students volunteer for a committee. How many different 3-person committees can be chosen?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

what do u do after u have 6?3(6-3)?????

OpenStudy (anonymous):

So you have 6 people, and you want to choose 3, so: \[\left(\begin{matrix}6 \\ 3\end{matrix}\right) = \frac{6!}{3!(6-3)!}\]

OpenStudy (anonymous):

\[=\frac{(6)(5)(4)(3)(2)(1)}{(3)(2)(1)(3)(2)(1)}\]

OpenStudy (anonymous):

what is that ? and what do i do with it ? do i divide or do i multiply???

OpenStudy (anonymous):

Simplifying by cancelling out the things that are the same in the numerator and denominator we get: \[\frac {(6)(5)(4)}{(3)(2)(1)} = \frac{120}{6} = 20\]

OpenStudy (anonymous):

oh i see now! ok let me save the formula .... 1 sec. DONT GO ANYWHERE! lol (please)

OpenStudy (anonymous):

You and two friends visit a pizza shop. There are 12 toppings. 7 toppings are meat and the rest are vegetable. Each of you orders a different topping (no repeats) at random. What is the probability that your group orders only meat toppings? i think the answer is 7/44 (im prob. wrong)

OpenStudy (anonymous):

Wow, alright let me think about this lol...you are going to pick 2 toppings. there are 12 total toppings, so: \[\left(\begin{matrix}12 \\ 2\end{matrix}\right) = 66\] is the total number of combinations (including vegies and all that). Now lets see how many of those combinations only have meat. That would be \[\left(\begin{matrix}7 \\ 2\end{matrix}\right) = 21\] So the probability is: \[\frac{21}{66} = \frac{7}{22}\]

OpenStudy (anonymous):

gtg

OpenStudy (anonymous):

ACK!! there are 3 people, my stuff is wrong, i just noticed, you are probably right

OpenStudy (anonymous):

alright, sry bout the mix up.

OpenStudy (anonymous):

Your answer is correct, just calculated it

OpenStudy (anonymous):

how does this make sense? O.o

OpenStudy (anonymous):

honestly i have NO CLUE! lol

OpenStudy (anonymous):

this is freshman stuff? i didnt go through this....